OpenStudy (anonymous):

The area of a triangle is 56.4 square inches and the base is 12 inches. What is the height of this triangle @luckycoins888

OpenStudy (anonymous):

A = 1/2 BH plug in what you have been provided and solve for the missing value (in this case H)

OpenStudy (anonymous):

maybe easier for you to rewrite the equation to solve for H..and then plug in

OpenStudy (anonymous):

? what

OpenStudy (danjs):

Area of Triangle = half of the base times the height

OpenStudy (danjs):

\[A = \frac{ 1 }{ 2 }*b*h\] 2*A = b*h

OpenStudy (danjs):

\[h = \frac{ 2*A }{ b }\] They give you Area A=56.4 and base b = 12 calculate height h

OpenStudy (anonymous):

?

OpenStudy (danjs):

Height equals two times the area divided by the base

OpenStudy (anonymous):

4.7

OpenStudy (danjs):

\[h = \frac{ 2 *A }{ b } = \frac{ 2 * 56.4 }{ 12 } = \]